# Exploring Textual Semantics Diversity for Image Transmission in Semantic Communication Systems using Visual Language Model

### v1

@@ V Conclusion
In this letter, we propose a VLM assisted multi-text semantic communication system by exploring multiple textual semantics, which greatly reduces the transmission amount of semantic features and realizes image transmission under low SNR conditions. We extracte the text and image features by VLM and FCN, respectively, and then use BART to correct spelling errors for reducing the impact of noise on the text. Finally, the image is reconstructed by combining the image features and texts to finish a complete image transmission process. Simulation results show that text features plus a small number of image features can complete the image transmission robustly, and a small number of image features can effectively limit the randomness of the generated images.

### v2

@@ V Conclusion
In this paper, we propose a VLM assisted multi-text semantic communication system by exploring multiple textual semantics, which greatly reduces the transmission amount of semantic features without reducing the effective extraction rate of features and realizes image transmission under low SNR conditions. We extract the image features by FCN and fuse the adjacent main features into one through the proposed fusion feature method. We extract the text features by VLM, and then use BART to correct spelling errors for reducing the impact of noise on the text. Finally, the image is reconstructed by combining the image features and texts to finish a complete image transmission process. The simulation results show that feature fusion can dramatically improve the accuracy of semantic features, and accurate text features with a small number of image features can robustly complete the image transmission, and a small number of image features can effectively limit the randomness of the generated images.